on monday harold picked five donuts and six large coffees he paid $8.65
on tuesday he picked three donuts and five coffees he paid $6.38. what's the cost of one donut and the cost of one large coffee

Answer by stanbon(75887)

on monday harold picked five donuts and six large coffees he paid $8.65
on tuesday he picked three donuts and five coffees he paid $6.38. what's the cost of one donut and the cost of one large coffee
------
Equation:
5d + 6c = 865 cents
3d + 5c = 638 cents
------------------------
Multiply thru 1st by 3:
Multiply thru 2nd by 5:
15d + 18c = 3*865
15d + 25c = 5*638
---
Subtract and solve for "c":
7c = 595
c = 75 cents (cost on one large coffee)
----
solve for "d":
3d + 5c = 638 cents
3d + 5*75 = 638
3d = 263
d = 87 2/3 cents (cost of one donut)
===============
Cheers,
Stan H.
===============